a passenger train and a freight train leave the same station at the same time. passenger train travels east at 60 mph and freight train travels west at 30 mph. in how many hours will they be 150 miles apart?
Answer by fcabanski(1391)

D = rt. Distance = rate times time.


t = Distance / rate.


Their rate distance increase is 60 + 30 = 90 mph. The distance between them increases by 90 miles every hour.


The desired distance is 150.


150/90 = 5/3 = 5 and 2/3 hours = 1 hour 40 minutes.